David is 3 years older than his brother. They have a combined age of 23. how old is David.
a. 7
b. 10
c. 13
d. 16

Answer by colliefan(242)

Write as an algebraic equation.
brother's age = x
David's age = x+3
The sum is 23 so:
x+(x+3)=23
2x+3=23
2x+3-3=23-3
2x+0=20
2x=20
2x*1/2=20*1/2
x*1=10
x=10
Check to see if 10 and 13 are solutions. 10+13 is 23.